# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ 4.1 Sliding-window probabilistic extraction procedure
To identify regions of memorization, we develop the following “panning for gold” approach. For a given book, we start at the beginning of the text file in Books3. We take a chunk of text that is sufficiently long to contain $100$ tokens, slide $10$ characters forward in the book text and repeat this process. We do this for the entire length of the book, which results in approximately $1$ sequence every $10$ characters, e.g., The Great Gatsby has 270,870270{,}870 characters and thus roughly 27,00027{,}000 sequences. For most of our experiments, we follow the standard procedure of evaluating $100$-token sequences [18, 60], which we divide into $50$-token prefix prompts and $50$-token suffixes that we attempt to extract (Equation 1).
@@ 4.2 Average extraction rates obscure book-specific memorization
Figure 4: Average extraction rates are low, but book-specific extraction varies widely. (left) Comparing extraction rates (Equation 3) of random $100$-token sequences ($50$-token prefix ++ $50$-token suffix) from Books3 for different models trained on Books3. We show the greedy (blue) and probabilistic (orange, $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$, top-$k$ decoding with T=1T{=}1 and k=40k{=}40) discoverable extraction rates (Appendix C). For both, average extraction is low. (right) Extraction coverage (Equation 4) for specific books reveals a more nuanced picture. For Harry Potter and the Sorcerer’s Stone [106] and Sandman Slim [69], we show coverage for $\tau\in\{1\%,10\%,50\%,75\%\}$ for $3$ models (Appendix D.4). Nearly half of Harry Potter and the Sorcerer’s Stone can be extracted with respect $p_{\bm{z}}\!\geq\!\tau\!=\!50\%$.
Our sliding-window procedure provides a detailed, work-specific view of memorization that is distinct from prior work in ML. That is, while prior work shows that models memorize certain amounts of the data they were trained on [e.g., 16, 93, 60, 74, 9, 48, 121, 8], it typically quantifies memorization through overall extraction rates—much like those in Figure 4 (left). For some (part of a) training dataset, researchers draw (typically at random) a set $\mathbb{Z}$} of sequences $\bm{z}$} of a specified length (e.g., $100$ tokens). For each sequence, they prompt with its prefix and count extraction as successful if the resulting generation matches the target suffix (Section 2). The extraction rate is computed as the number of attempted extractions that succeeded, relative to the total number of attempts:
@@
These reported averages are generally small, just as we observe in Figure 4 (left) for $9$ models trained on Books3. The highest rate we observe on random $100$-token sequences ($50$-token prefix ++ $50$-token suffix) from Books3 is just over $0.6\%$ (for Llama 3.1 70B). Small average extraction rates often form the basis of defendants in copyright infringement suits calling memorization a rare “bug” [28, 97]. However, while average extraction rates are useful for estimating overall memorization of Books3 text, they clearly tell an incomplete story. Low extraction rates signal that models likely do not memorize most text in Books3, but specific works can be highly memorized. On average, Llama 3.1 70B may have only memorized about $0.6\%$ of sequences in Books3, but concealed within that small number is that it memorized effectively all of 1984 (Figure 3).
We can see this intuitively in Figure 3; target suffixes whose extraction probability exceeds $\tau_{\text{min}}{=}0.1\%$ can be found nearly everywhere in the book. [^5] Exceptions are this particular edition’s forward at the beginning of the book and notes near the end of the book—located before some selections of other writing by George Orwell, which are also highly memorized. However, we can also make this more concrete: using the results of our sliding-window procedure, we can quantify how much of a whole book is extractable. We compute extraction coverage as the fraction of characters in a book that lie within at least one suffix whose extraction probability $p_{\bm{z}}$ exceeds a threshold $\tau$ (Appendix D.4)—where $\tau\in[\tau_{\text{min}},1]$ because $\tau_{\text{min}}$ is the minimum probability we consider to reflect extraction success (Sections 2.15). Let $\mathsf{span}({\bm{z}})$ be the set of character indices spanned by a given $50$-token suffix of sequence $\bm{z}$}. Then
@@
For a more detailed comparison of book-specific coverage, in Figure 4 (right) we show extraction coverage for Harry Potter and the Sorcerer’s Stone [106] and Sandman Slim [69]—one of the books by plaintiff Richard Kadrey in Kadrey et al. v. Meta, Inc. [70]). Sandman Slim is hardly memorized at all. Extraction coverage for this book is less than $0.4\%$ for all models and thresholds $\tau$. [^7] Prior studies (and Figure 4) show that larger LLMs exhibit higher extraction rates [18]. This pattern does not cleanly generalize to specific works. At $\tau{=}75\%$, Llama 3.1 70B exhibits $0.00\%$ extraction coverage of Sandman Slim, while the smaller Llama 1 65B and Pythia 12B exhibit $0.09\%$ and $0.22\%$, respectively. In contrast, the first Harry Potter [106] is highly memorized in Llama models (Section 6). For Llama 3.1 70B, $43.26\%$ of the book is covered with $\tau\!=\!50\%$.888 The version we evaluate from Book3 includes the first chapter of the second Harry Potter book at the end. For lower $\tau$, there is higher coverage; $90.89\%$ of the book is covered with $\tau\!=\!1\%$. While lower thresholds correspond to extracting text less reliably—i.e., requiring more prompts to the model—for an LLM, they reflect enormous probabilities (Section 2.1).
@@ 5 Validating our extraction procedure
Figure 6: Evaluating Llama 3.1 70B, Llama 3.1 8B, and Phi 4 on We Were Eight Years in Power [23] (a book in Books3), using prefix lengths in $\{25,50,100,200,400,800\}$ and $50$-token suffixes with top-$k$ decoding (T=1T{=}1, k=40k{=}40). The book is in Llama’s training data, but not Phi 4’s. Longer prefixes reveal more memorized sequences, though not all sequences in the book are extractable. For Phi 4, the only extractable suffixes are widely duplicated quotes that likely appeared in its training data from other sources. These results support that when our procedure identifies extraction, it is detecting true instances of memorization rather than chance generations.
As a baseline, we run experiments on a book from Books3 and Llama 3.1 models, which were trained on Books3 (Section 5.1). As expected [18], longer prefixes surface more memorized training data. Then, we run a set of negative controls: we use the same procedure for models on non-training data, which tests whether our method spuriously registers false positives on data where memorization is impossible (Section 5.2). Unlike the experiments on training data, regardless of prefix length, suffix probabilities remain below $\tau_{\text{min}}$ (with interesting exceptions [^9] These sequences are boilerplate text (copyright notices, publisher addresses, etc.) and popular quotes—both of which are duplicated across many sources aside from Books3. (See Appendices D.5E.6, & G.) ), demonstrating our approach does not mistakenly register non-training data as extractable. Together, these experiments show that, when our measurements indicate extraction, we can be confident they are a valid reflection of memorization. We briefly discuss the implications of these results (Section 5.3), which inform the experiments we run in the following section. Additional results and discussion can be found in Appendix E.
@@ 5.1 Longer prefixes surface more memorized training data
When a sequence is known to be in the training data, memorization is the most plausible explanation for high-probability generation of its (sufficiently long) verbatim suffix given the corresponding prefix (Section 2). So, as a baseline, we apply the sliding-window procedure to this setting (Figure 6): we attempt to extract sequences in We Were Eight Years in Power [23] (a book in Books3) from Llama 3.1 70B and Llama 3.1 8B (models trained on Books3). Longer prefixes reveal additional memorized sequences. This is what Carlini et al. [18] call the discoverability phenomenon: providing additional context (a longer ground-truth prefix) can raise the conditional probability of the target suffix $p_{\bm{z}}$ enough that it becomes extractable ($p_{\bm{z}}\!\geq\!\tau_{\text{min}}$). However, many sequences remain unextractable: no matter how long the prefix, their suffix probabilities $p_{\bm{z}}$ never surpass $\tau_{\text{min}}$. [^10] Our results suggest that longer prefixes beyond $800$ tokens would enable extraction of additional sequences from this book, but the gains diminish, indicating that not every sequence is extractable (Appendix E.3).
For example, Llama 3.1 70B memorizes most of the book, yet consistent sections remain unextractable across all prefix lengths (e.g., white gaps in Figure 6, such as from ${\sim}75$K–120120K characters). The same is even clearer for Llama 3.1 8B: being smaller, it memorizes significantly less [18, 60, 16]. Longer prefixes do surface more memorized training data, but most of the book remains unextractable. These results are consistent with prior work: high-quality models do not—indeed cannot [74, 16, 84, 26]—memorize all of the training data in their massive training datasets. Even as prefix length increases, our measurements do not suddenly capture extraction success for everything; they capture success for sequences that have been memorized. [^11] We observe the same for Llama 3.1 70B on other books in its training data. Some appear effectively not memorized at all, with only limited increases in extraction success at longer prefix lengths (Appendix E.3). We further support this point in the next section, where we compare these results to experiments on non-training data.
@@ 5.2 Measurements on non-training data do not mistakenly register extraction
We run two sets of negative controls. First, we apply our procedure to Books3 books with Phi 4 [2]—a model that was not trained on whole copyrighted books. Second, we run Llama 3.1 70B (trained on Books3), on books from 2025, i.e., books that post-date its 2024 training cutoff. In our plots, we distinguish these conditions by color: blue for models trained on Books3 on Books3 text, orange for Phi 4 on Books3 text, and gray for results on post-cutoff books.
@@
Negative control 1: A model not trained on Books3 $\times$ Books3 text.  We run the same experiment as in Section 5.1 on Phi 4, which was not trained on in-copyright books like We Were Eight Years in Power [23]. To support the validity of our measurement approach, unlike with Llama 3.1 models, we should not observe extraction success for unique text from this book. Our results for Phi 4 in Figure 6 confirm this expectation: increasing the context length does not push the conditional probability of non-training-data suffixes high enough to cross the $\tau_{\text{min}}$ threshold. We do observe increased extraction signal, but it only concentrates in four small areas that do not appear to be false positives. Instead, they plausibly reflect memorized text: the suffixes our procedure counts as extracted are famous quotes that are widely duplicated on the Internet, and thus very likely to have been included in Phi 4’s training data via sources other than this copyrighted book [2]. For example, at ${\sim}280$K characters, we extract sequences containing passages from the King James Bible [124]; at ${\sim}546$K characters, we extract a quote from Barack Obama’s 2004 Democratic National Convention speech [95]. Notably, Llama 3.1 models also exhibit high extraction probabilities in these same regions. While they were trained on this book, it is plausible that duplicates from other sources contributed to the memorization and extractability of these sequences [74].
@@
Figure 7: Validating probabilistic extraction by computing $p_{\bm{z}}$ for non-training data: Llama 3.1 70B and top-$k$ (T=1T{=}1, k=40k{=}40) decoding on Great Big Beautiful Life [63] (published spring 2025), with respect to $50$-token suffixes and prefix lengths in $\{200,400,800\}$ tokens. Except for the copyright notice at the beginning, we do not register any $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$. We register no other suffix probabilities $p_{\bm{z}}\!\geq\!10^{-13}\%$, and only 11 (out of ${\sim}62{,}000$ sequences) for which $p_{\bm{z}}{\geq}10^{-14}\%$.
@@
Negative control 2: Models trained on Books3 $\times$ books published after their training.  We apply the same procedure to Llama models on books published in 2025—after the July 2024 cutoff for Llama 3.1 training. In Figure 7, we show one example for Great Big Beautiful Life [63]. Apart from the (highly duplicated, boilerplate) copyright notice, for all prefix lengths there are no suffix probabilities that, in analogous experiments on training data, we would count as successful extraction. Not only does $p_{\bm{z}}$ never surpass $\tau_{\text{min}}\!=\!0.1\%\!=\!10^{-1}\%$, there are no sequences with $p_{\bm{z}}\geq 10^{-13}\%$. Exactly $1$ sequence (out of ${\sim}62{,}000$) has $p_{\bm{z}}\geq 10^{-14}\%$ (the faint gray band at ${\sim}20$K characters). This reflects chance generation of non-training data—effectively a “monkey at the typewriter” event. But importantly, this is not a sequence that our method counts as extractable: the suffix probability is $13$ orders of magnitude smaller than our setting for $\tau_{\text{min}}$. With a single attempt, the odds of generating the verbatim suffix are $\nicefrac{{1}}{{p_{\bm{z}}}}\!\approx\!\nicefrac{{1}}{{10^{-14}\%}}$, or $1$ in $10$ quadrillion (Appendix E.4.2). To give a sense of just how small this number is, the chance of winning the Powerball jackpot in a single draw is about $1$ in $292$ million—roughly $34$ million times more likely.

### v4

@@ 4.1 Sliding-window probabilistic extraction procedure
To identify regions of memorization, we develop the following “panning for gold” approach. For a given book, we start at the beginning of the text file in Books3. We take a chunk of text that is sufficiently long to contain $100$ tokens, slide $10$ characters forward in the book text and repeat this process. We do this for the entire length of the book, which results in approximately $1$ sequence every $10$ characters, e.g., The Great Gatsby has 270,870270{,}870 characters and thus roughly 27,00027{,}000 sequences. For most of our experiments, we follow the standard procedure of evaluating $100$-token sequences [carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels], which we divide into $50$-token prefix prompts and $50$-token suffixes that we attempt to extract (Equation 1).
@@ 4.2 Average extraction rates obscure book-specific memorization
Figure 4: Average extraction rates are low, but book-specific extraction varies widely. (left) Comparing extraction rates (Equation 3) of random $100$-token sequences ($50$-token prefix ++ $50$-token suffix) from Books3 for different models trained on Books3. We show the greedy (blue) and probabilistic (orange, $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$, top-$k$ decoding with T=1T{=}1 and k=40k{=}40) discoverable extraction rates (Appendix C). For both, average extraction is low. (right) Extraction coverage (Equation 4) for specific books reveals a more nuanced picture. For Harry Potter and the Sorcerer’s Stone [hp1] and Sandman Slim [sandmanslim], we show coverage for $\tau\in\{1\%,10\%,50\%,75\%\}$ for $3$ models (Appendix D.4). Nearly half of Harry Potter and the Sorcerer’s Stone can be extracted with respect $p_{\bm{z}}\!\geq\!\tau\!=\!50\%$.
Our sliding-window procedure provides a detailed, work-specific view of memorization that is distinct from prior work in ML. That is, while prior work shows that models memorize certain amounts of the data they were trained on [e.g., carlini2021extracting, nasr2023scalable, hayes2025measuringmemorizationlanguagemodels, lee2022dedup, biderman2023pythia, reid2024gemini, gemma2, biderman2023emergent], it typically quantifies memorization through overall extraction rates—much like those in Figure 4 (left). For some (part of a) training dataset, researchers draw (typically at random) a set $\mathbb{Z}$} of sequences $\bm{z}$} of a specified length (e.g., $100$ tokens). For each sequence, they prompt with its prefix and count extraction as successful if the resulting generation matches the target suffix (Section 2). The extraction rate is computed as the number of attempted extractions that succeeded, relative to the total number of attempts:
@@
These reported averages are generally small, just as we observe in Figure 4 (left) for $9$ models trained on Books3. The highest rate we observe on random $100$-token sequences ($50$-token prefix ++ $50$-token suffix) from Books3 is just over $0.6\%$ (for Llama 3.1 70B). Small average extraction rates often form the basis of defendants in copyright infringement suits calling memorization a rare “bug” [cooper2024files, openairesponse]. However, while average extraction rates are useful for estimating overall memorization of Books3 text, they clearly tell an incomplete story. Low extraction rates signal that models likely do not memorize most text in Books3, but specific works can be highly memorized. On average, Llama 3.1 70B may have only memorized about $0.6\%$ of sequences in Books3, but concealed within that small number is that it memorized effectively all of 1984 (Figure 3).
We can see this intuitively in Figure 3; target suffixes whose extraction probability exceeds $\tau_{\text{min}}{=}0.1\%$ can be found nearly everywhere in the book. [^5] Exceptions are this particular edition’s forward at the beginning of the book and notes near the end of the book—located before some selections of other writing by George Orwell, which are also highly memorized. However, we can also make this more concrete: using the results of our sliding-window procedure, we can quantify how much of a whole book is extractable. We compute extraction coverage as the fraction of characters in a book that lie within at least one suffix whose extraction probability $p_{\bm{z}}$ exceeds a threshold $\tau$ (Appendix D.4)—where $\tau\in[\tau_{\text{min}},1]$ because $\tau_{\text{min}}$ is the minimum probability we consider to reflect extraction success (Sections 2.1 & 5). Let $\mathsf{span}({\bm{z}})$ be the set of character indices spanned by a given $50$-token suffix of sequence $\bm{z}$}. Then
@@
For a more detailed comparison of book-specific coverage, in Figure 4 (right) we show extraction coverage for Harry Potter and the Sorcerer’s Stone [hp1] and Sandman Slim [sandmanslim]—one of the books by plaintiff Richard Kadrey in Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated]). Sandman Slim is hardly memorized at all. Extraction coverage for this book is less than $0.4\%$ for all models and thresholds $\tau$. [^7] Prior studies (and Figure 4) show that larger LLMs exhibit higher extraction rates [carlini2023quantifying]. This pattern does not cleanly generalize to specific works. At $\tau{=}75\%$, Llama 3.1 70B exhibits $0.00\%$ extraction coverage of Sandman Slim, while the smaller Llama 1 65B and Pythia 12B exhibit $0.09\%$ and $0.22\%$, respectively. In contrast, the first Harry Potter [hp1] is highly memorized in Llama models (Section 6). For Llama 3.1 70B, $43.26\%$ of the book is covered with $\tau\!=\!50\%$. [^8] The version we evaluate from Book3 includes the first chapter of the second Harry Potter book at the end. For lower $\tau$, there is higher coverage; $90.89\%$ of the book is covered with $\tau\!=\!1\%$. While lower thresholds correspond to extracting text less reliably—i.e., requiring more prompts to the model—for an LLM, they reflect enormous probabilities (Section 2.1).
@@ 5 Validating our extraction procedure
Figure 6: Evaluating Llama 3.1 70B, Llama 3.1 8B, and Phi 4 on We Were Eight Years in Power [wewere] (a book in Books3), using prefix lengths in $\{25,50,100,200,400,800\}$ and $50$-token suffixes with top-$k$ decoding (T=1T{=}1, k=40k{=}40). The book is in Llama’s training data, but not Phi 4’s. Longer prefixes reveal more memorized sequences, though not all sequences in the book are extractable. For Phi 4, the only extractable suffixes are widely duplicated quotes that likely appeared in its training data from other sources. These results support that when our procedure identifies extraction, it is detecting true instances of memorization rather than chance generations.
As a baseline, we run experiments on a book from Books3 and Llama 3.1 models, which were trained on Books3 (Section 5.1). As expected [carlini2023quantifying], longer prefixes surface more memorized training data. Then, we run a set of negative controls: we use the same procedure for models on non-training data, which tests whether our method spuriously registers false positives on data where memorization is impossible (Section 5.2). Unlike the experiments on training data, regardless of prefix length, suffix probabilities remain below $\tau_{\text{min}}$ (with interesting exceptions [^9] These sequences are boilerplate text (copyright notices, publisher addresses, etc.) and popular quotes—both of which are duplicated across many sources aside from Books3. (See Appendices D.5, E.6, & G.) ), demonstrating our approach does not mistakenly register non-training data as extractable. Together, these experiments show that, when our measurements indicate extraction, we can be confident they are a valid reflection of memorization. We briefly discuss the implications of these results (Section 5.3), which inform the experiments we run in the following section. Additional results and discussion can be found in Appendix E.
@@ 5.1 Longer prefixes surface more memorized training data
When a sequence is known to be in the training data, memorization is the most plausible explanation for high-probability generation of its (sufficiently long) verbatim suffix given the corresponding prefix (Section 2). So, as a baseline, we apply the sliding-window procedure to this setting (Figure 6): we attempt to extract sequences in We Were Eight Years in Power [wewere] (a book in Books3) from Llama 3.1 70B and Llama 3.1 8B (models trained on Books3). Longer prefixes reveal additional memorized sequences. This is what carlini2023quantifying call the discoverability phenomenon: providing additional context (a longer ground-truth prefix) can raise the conditional probability of the target suffix $p_{\bm{z}}$ enough that it becomes extractable ($p_{\bm{z}}\!\geq\!\tau_{\text{min}}$). However, many sequences remain unextractable: no matter how long the prefix, their suffix probabilities $p_{\bm{z}}$ never surpass $\tau_{\text{min}}$. [^10] Our results suggest that longer prefixes beyond $800$ tokens would enable extraction of additional sequences from this book, but the gains diminish, indicating that not every sequence is extractable (Appendix E.3).
For example, Llama 3.1 70B memorizes most of the book, yet consistent sections remain unextractable across all prefix lengths (e.g., white gaps in Figure 6, such as from ${\sim}75$K–120120K characters). The same is even clearer for Llama 3.1 8B: being smaller, it memorizes significantly less [carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels, carlini2021extracting]. Longer prefixes do surface more memorized training data, but most of the book remains unextractable. These results are consistent with prior work: high-quality models do not—indeed cannot [lee2022dedup, carlini2021extracting, mahdavi2024memorizationcapacitymultiheadattention, collins2017capacitytrainabilityrecurrentneural]—memorize all of the training data in their massive training datasets. Even as prefix length increases, our measurements do not suddenly capture extraction success for everything; they capture success for sequences that have been memorized. [^11] We observe the same for Llama 3.1 70B on other books in its training data. Some appear effectively not memorized at all, with only limited increases in extraction success at longer prefix lengths (Appendix E.3). We further support this point in the next section, where we compare these results to experiments on non-training data.
@@ 5.2 Measurements on non-training data do not mistakenly register extraction
We run two sets of negative controls. First, we apply our procedure to Books3 books with Phi 4 [phi4]—a model that was not trained on whole copyrighted books. Second, we run Llama 3.1 70B (trained on Books3), on books from 2025, i.e., books that post-date its 2024 training cutoff. In our plots, we distinguish these conditions by color: blue for models trained on Books3 on Books3 text, orange for Phi 4 on Books3 text, and gray for results on post-cutoff books.
@@
Negative control 1: A model not trained on Books3 $\times$ Books3 text.  We run the same experiment as in Section 5.1 on Phi 4, which was not trained on in-copyright books like We Were Eight Years in Power [wewere]. To support the validity of our measurement approach, unlike with Llama 3.1 models, we should not observe extraction success for unique text from this book. Our results for Phi 4 in Figure 6 confirm this expectation: increasing the context length does not push the conditional probability of non-training-data suffixes high enough to cross the $\tau_{\text{min}}$ threshold. We do observe increased extraction signal, but it only concentrates in four small areas that do not appear to be false positives. Instead, they plausibly reflect memorized text: the suffixes our procedure counts as extracted are famous quotes that are widely duplicated on the Internet, and thus very likely to have been included in Phi 4’s training data via sources other than this copyrighted book [phi4]. For example, at ${\sim}280$K characters, we extract sequences containing passages from the King James Bible [deuteronomy]; at ${\sim}546$K characters, we extract a quote from Barack Obama’s 2004 Democratic National Convention speech [obama2004]. Notably, Llama 3.1 models also exhibit high extraction probabilities in these same regions. While they were trained on this book, it is plausible that duplicates from other sources contributed to the memorization and extractability of these sequences [lee2022dedup].
@@
Figure 7: Validating probabilistic extraction by computing $p_{\bm{z}}$ for non-training data: Llama 3.1 70B and top-$k$ (T=1T{=}1, k=40k{=}40) decoding on Great Big Beautiful Life [greatbig] (published spring 2025), with respect to $50$-token suffixes and prefix lengths in $\{200,400,800\}$ tokens. Except for the copyright notice at the beginning, we do not register any $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$. We register no other suffix probabilities $p_{\bm{z}}\!\geq\!10^{-13}\%$, and only 11 (out of ${\sim}62{,}000$ sequences) for which $p_{\bm{z}}{\geq}10^{-14}\%$.
@@
Negative control 2: Models trained on Books3 $\times$ books published after their training.  We apply the same procedure to Llama models on books published in 2025—after the July 2024 cutoff for Llama 3.1 training. In Figure 7, we show one example for Great Big Beautiful Life [greatbig]. Apart from the (highly duplicated, boilerplate) copyright notice, for all prefix lengths there are no suffix probabilities that, in analogous experiments on training data, we would count as successful extraction. Not only does $p_{\bm{z}}$ never surpass $\tau_{\text{min}}\!=\!0.1\%\!=\!10^{-1}\%$, there are no sequences with $p_{\bm{z}}\geq 10^{-13}\%$. Exactly $1$ sequence (out of ${\sim}62{,}000$) has $p_{\bm{z}}\geq 10^{-14}\%$ (the faint gray band at ${\sim}20$K characters). This reflects chance generation of non-training data—effectively a “monkey at the typewriter” event. But importantly, this is not a sequence that our method counts as extractable: the suffix probability is $13$ orders of magnitude smaller than our setting for $\tau_{\text{min}}$. With a single attempt, the odds of generating the verbatim suffix are $\nicefrac{{1}}{{p_{\bm{z}}}}\!\approx\!\nicefrac{{1}}{{10^{-14}\%}}$, or $1$ in $10$ quadrillion (Appendix E.4.2). To give a sense of just how small this number is, the chance of winning the Powerball jackpot in a single draw is about $1$ in $292$ million—roughly $34$ million times more likely.